OpenStudy (anonymous):

consider the function f(x)=3x-8 A.) find the inverse of f(x) and name it g(x). B.) use the function composition to show that f(x) and g(x) are inverses of each other. C.) draw a graph of f(x) and g(x).

OpenStudy (anonymous):

i already have a down i just need help with b and c

OpenStudy (anonymous):

@rockstar0765

OpenStudy (anonymous):

@chosenmatt

OpenStudy (rockstar0765):

the inverse like as in what

OpenStudy (anonymous):

inverse of function f(X) its like the opposite i think its g(x)= (x+8)/3

OpenStudy (anonymous):

Let's write y=f(x) Let g be the inverse of f. g is inverse of if and only if x=g(y) we have: f(x) =3x-8 or y=3x-8 y+8=3x x=(y+8)/3 The inverse function of f(x) is g(y) = (Y+8)/3 OR g(y)=y/3+8/3 OR g(y)=(1/3)y +8/3

OpenStudy (anonymous):

thank you so much but I already got it

OpenStudy (anonymous):

You're welcome. I saw your answer, but, sorry! I wanted to call your attention a little bit on a mistake. The inverse function should be a function of y, not a function of x. I mean the variable is no more x, but y.